how long will it take you to triple an amount of $500 if you invest it at a rate of 6% compounded annually?
Answer by Boreal(15235)

P=Po(1+.06)^n
1500=500(1.06)^n
3=1.06^n
ln3=n ln 1.06
n=1n3/ln1.06=18.85 years
Doubling is the rule of 70 (or 72, since 72 is evenly divisible by more integers)
Tripling is the rule of 110, which is ln 3*100 to use the percentage of interest rate.
Using that rule, you can solve such problems by dividing 6 into 110, which would be 18.3 years, very close to the above answer.